### v2

# GP-MoLFormer-Sim: Test Time Molecular Optimization through Contextual Similarity Guidance

Abstract
The ability to design molecules while preserving similarity to a target molecule and/or property is crucial for various applications in drug discovery, chemical design, and biology. We introduce in this paper an efficient training-free method for navigating and sampling from the molecular space with a generative Chemical Language Model (CLM), while using the molecular similarity to the target as a guide. Our method leverages the contextual representations learned from the CLM itself to estimate the molecular similarity, which is then used to adjust the autoregressive sampling strategy of the CLM. At each step of the decoding process, the method tracks the distance of the current generations from the target and updates the logits to encourage the preservation of similarity in generations. We implement the method using a recently proposed $\sim$47M parameter SMILES-based CLM, GP-MoLFormer, and therefore refer to the method as GP-MoLFormer-Sim, which enables a test-time update of the deep generative policy to reflect the contextual similarity to a set of guide molecules. The method is further integrated into a genetic algorithm (GA) and tested on a set of standard molecular optimization benchmarks involving property optimization, molecular rediscovery, and structure-based drug design. Results show that, GP-MoLFormer-Sim, combined with GA (GP-MoLFormer-Sim+GA) outperforms existing training-free baseline methods, when the oracle remains black-box. The findings in this work are a step forward in understanding and guiding the generative mechanisms of CLMs.
1 Introduction
Finding new functional molecules with desired structure and properties involves solving a constrained multi-objective optimization problem, which is crucial in many applications such as drug discovery and new material design. Given the large size of the molecular space, brute-force search around known substructures is often inefficient and costly for such tasks. Existing molecular optimization algorithms therefore mainly involve reinforcement learning, deep generative models, genetic algorithms, or a combination thereof. Recent works show that traditional genetic algorithm (GA)-based methods with domain-specific operators are competitive when compared to costlier alternatives that involve deep learning models gao2022sample; tripp2023geneticalgorithmsstrongbaselines. Earlier efforts that have successfully combined GAs with deep learning for better search typically require further training of the deep learning model, more specifically of the deep generative model, to adapt the generative policy for generating high-reward samples corresponding to the specific optimization problem ahn2020guiding; kim2024geneticguidedgflownetssampleefficient.
Different from earlier approaches, here we propose a training-free method for equipping a pre-trained deep generative model for targeted search. The proposed method exploits the contextual similarity between a target molecule and a set of generated molecules with a generative chemical language model (CLM). We update the autoregressive decoding policy of the generative model on the fly as a means to guide the generation toward high-reward samples. We use the recently proposed SMILES-based GP-MoLFormer model ross2025gpmolformerfoundationmodelmolecular as the base generative CLM to generate molecules, and therefore refer to this test-time contextual similarity-based guided generation method as GP-MoLFormer-Sim. We first show the performance of GP-MoLFormer-Sim on a similarity-based lead optimization task where the goal is to generate molecules of high similarity with respect to a given target molecule in a sample-efficient manner. Experiments on this task show that the proposed method outperforms random search as well as a reinforcement learning-based baseline.
We further integrate GP-MoLFormer-Sim with a genetic algorithm-based search process, where GP-MoLFormer-Sim enables generating offspring of the high-reward samples (see Figure 1). We refer to this approach as GP-MoLFormer-Sim+GA. Results on the popular Practical Molecular Optimization (PMO) benchmark gao2022sample show that GP-MoLFormer-Sim+GA yields better performance on 23 molecular optimization tasks when the oracle is a black-box, compared to the current state-of-the-art GA-based training-free baselines including the ones that call large language models like GPT-4 for proposing high-reward samples. To our knowledge, this is the first demonstration of using test-time update of a CLM-based deep generative policy for for molecular optimization.
2 Guided Generation
Figure 1: Overview of the GP-MoLFormer-Sim+GA process. (A) The top-k highest scoring molecules so far are chosen as guides with additional diverse candidates, if desired. (B) Using GP-MoLFormer-Sim, generate new candidates conditioned on closeness to top guides. GP-MoLFormer-Sim adjusts the logits of the base model at every iteration using embedding similarity of the guide sequence so far to each proposed next token. (C) Prune (filter) or augment (with graph-based crossover operation) generations and score them with the oracle. These new samples are added back to the population and the process is repeated until the oracle budget is met.
2.1 Background information — GP-MoLFormer
GP-MoLFormer is a chemical language foundation model, which is a GPT-style autoregressive decoder trained with linear attention and rotary embeddings ross2025gpmolformerfoundationmodelmolecular [^1] Available via https://huggingface.co/ibm-research/GP-MoLFormer-Uniq. The model used in here is trained on $\sim$650M canonicalized SMILES obtained from ZINC and PubChem databases. Unconditional sampling from this chemical language model would allow exploring the chemical space. For details of the GP-MoLFormer model and its performance on unconditional SMILES generation task, see ross2025gpmolformerfoundationmodelmolecular.
2.2 Target-guided generation with $GP-MoLFormer$-Sim
Guiding the autoregressive sampling from CLMs like GP-MoLFormer towards specific molecules is of a paramount interest, as it enables generating new variations given target molecules of importance. Specifically, given a single molecule we are interested in exploring the molecular neighborhood where the similarity is defined through the cosine in the embedding space of the same generative model (GP-MoLFormer).
[…]
1:Inputs: $\alpha$ (mixing parameter), $\tau$ (softmax temperature), $T$ (RFF)
[…]
4:while EOS is not met do
5:  Append and Embed Generated
6:  for $i\in\text{Vocab}$ do
7:   $x_{i}=\textsc{GetEmbeddingGPT}(s\oplus i)$
8:   $x_{i}\leftarrow RandomFeatures(x_{i})$      // Optional
[…]
10:  end for
11:  Embed targeted molecule up to time $t$
12:  for $j\in\text{targetMolecules}$ do
13:   $y_{j}=\textsc{GetEmbeddingGPT}(m_{j}[1:t])$
14:   $y_{j}\leftarrow RandomFeatures(y_{j})$      // Optional
[…]
16:  end for
17:  Compute pairwise cosine $(\text{Vocabsize}\times N$ where $N=\text{number of targetMolecules})$
18:  $S_{ij}=\langle x_{i},y_{j}\rangle,i=1\dots\text{Vocabsize}$, $j=1\dots N$
19:  $\bar{S}_{i}=\frac{1}{N}\sum_{j=1}^{N}S_{ij}$, for $i=1\dots\text{Vocabsize}$
20:  Tilting the logits
21:  $u\leftarrow\textsc{GetLogits}(s)$ (vector of size Vocabsize, if topk used this is k)
22:  Standardize $u$ and $\bar{S}$
23:  $u\leftarrow\frac{1}{\tau}((1-\alpha)u+\alpha\bar{S})$
24:  Sample with probability $\textsc{softmax}(u)$ and get token $d$
25:  $s\leftarrow s\oplus d$
26:  $t\leftarrow t+1$
27:end while
28:return $s$
Algorithm 1 Guided Generation with kernel approximation
More formally, we wish to generate a new sequence $s$ with guidance from molecules $m_{j},j=1\dots N$ that are canonical SMILES sequences. We build the sequence $s$ incrementally by gradually sampling tokens from a new policy that mixes the likelihood under GP-MoLFormer (logit $u$) and the contextual similarity of the new sequence to target molecules in the embedding space of GP-MoLFormer, $\bar{S}$. Algorithm 1 summarizes this procedure. The logits of the new guiding policy are $\frac{1}{\tau}((1-\alpha)u+\alpha\bar{S})$, where $\alpha\in[0,1]$ controls the mixing strength, interpolating between unconditional sampling from GP-MoLFormer and pure similarity based sampling, and $\tau$ is a sampling temperature that allows control over the entropy of the sampling. The sequence is generated iteratively until the <eos> token is selected.
Note that the cosine similarity is used to “tilt” the logits of the GP-MoLFormer using similarity to a neighborhood formed by target molecules in the embedding space. We can push this idea further using a kernel density estimator (KDE) with a gaussian kernel. The temperature $T$ of the kernel induces further locality control. We can approximate the KDE using Random Fourier Features (rff) (lines 8 and 14 in Algorithm 1).
Because our proposed algorithm is a test time guidance algorithm, it does not need any training procedure and enjoys multiple advantages of scalability, parallelism, efficiency, and versatility in its applicability to multiple domains. The complexity of the algorithm is linear in the vocabulary size (2362), the dimensionality of GP-MoLFormer embedding (768) and the number of target molecules. For experiments with random features, we also used $768$ random features. As our method is training-free, we do not require any computational resources or timing observations for training. A single A100 GPU is used for each inference task (we also note that GPU memory is not a concern as our memory footprint only occupied no more than 8 GB of VRAM). Even though our guided method has many more moving parts and occupies much more memory than the base model unconditional generation, the extra computation cost of the guided method is only roughly four times slower than the unconditional generation. As an example, generating a single token from the unconditional model takes, on average, $0.013$ seconds while generating a guided token takes $0.049$ seconds. When generating a batch of 20 molecules, the runtime is $1.02$ seconds (producing avg. molecule length of 43 tokens) for unconditional and $3.97$ seconds for guided generation (of molecules with 40 tokens on average).
In Figure 2, we showcase a depiction of the Algorithm 1 in action on a guided generation task (for details of the task, see Section 4), which aims to generate molecules within the individual neighborhood of five trypsin inhibitor targets (large circles) with varying guiding strength $\alpha$ and sampling temperature $\tau$ settings. In blue, we see the unconditional generation from GP-MoLFormer. When comparing the first two panels for the same guidance strength $\alpha$, we see the effect of sampling temperature $\tau$: the higher temperature ($\tau=0.40$) leads to a higher entropy resulting in a larger spread around the target molecules. On the other hand, comparing the outer two panels, for fixed sampling temperature $\tau=0.20$, we observe that larger mixing $\alpha=0.5$ leads to tighter clustering around the target molecules, away from the unconditional baseline in blue. It should be mentioned that the algorithm is not specific to using a single guide, and can be extended to guiding the generation to multiple target molecules simultaneously. Additional visualizations involving a guidance by multiple targets simultaneously can be found in the Appendix Figure 5.
Theoretical Justification
It can be shown that the computations on lines 23–24 in Algorithm 1 represent a solution of a closed-form optimization problem involving token similarity to the target molecule distribution and a KL divergence from the smoothed GP-MoLFormer distribution. Full detail is laid out in Appendix A.
Figure 2: Guided generations (dots) around five trypsin inhibitor targets (large circles) with various guiding strength $\alpha$ and sampling temperature $\tau$ settings. “Base” molecules from un-guided generation. Green, orange, red, purple and brown correspond to five different targets. Molecules are visualized in GP-MoLFormer embeddings space projected onto two t-SNE dimensions.
2.3 GP-MoLFormer-Sim augmented with genetic algorithm — GP-MoLFormer-Sim+GA
Algorithm 2 GP-MoLFormer-Sim+GA
0: Oracle $F:\mathbb{M}\rightarrow\mathbb{R}$, GP-MoLFormer-Sim()
1: Inputs: $G,K,B,D$
2: generation $\leftarrow$ 0, budget $\leftarrow$ 0
3: Initialize $P\sim\mathrm{ZINC}$
4: Store scores $R[P]\leftarrow F(P)$
5: Record (budget spent, generation, avg. $K$ best scores)
6:while Oracle budget $\leq$ B do
7:  Sample $S\subset P$      // Select $G$ best candidates as guides
8:  Optional: $S\leftarrow S\cup(S^{\prime}\subset P)$      // Augment by $D$ diverse candidates
9:  Generate $N\leftarrow\textsc{GP-MoLFormer-Sim}(s)\quad\forall s\in S$ // Use GP-MoLFormer-Sim to create neighbors (mutations) of guides
10:  Select $P^{\prime}\subset P\cup N$      // e.g, best by $T_{sim}$ to current top in $P$
11:  Optional: augment $P^{\prime}$      // e.g., via “graph-based” crossover of current best guides
12:  generation $\leftarrow$ generation + 1
13:  Store $R[P^{\prime}\setminus P]\leftarrow F(P^{\prime}\setminus P)$
14:  Set $P\leftarrow P^{\prime}$
15:  Record (budget spent, generation, avg. $K$ best scores)
16:end while
17:return Array of tuples (generation, budget spent, avg. top-$K$ score)
[…]
3 Related Work
Molecule optimization The goal of molecule optimization is to iteratively modify molecule structures to improve desired properties like binding affinity, solubility, drug likeliness, etc. The ability to represent molecules using text-based encodings like SMILES weininger1988smiles and SELFIES Krenn_2020_selfies, enables the application of natural language processing techniques to tackle this problem. Existing methods have used techniques such as reinforcement learning (RL) olivecrona2017molecular; blaschke2018application; loeffler2024reinvent; segler2017generating; neil2018exploring; zhou2019optimization, variational autoencoders gomez2018automatic; jin2018junction, Bayesian optimization moss2020boss; tripp2021fresh, GFlowNets shen2023tacogfn; bengio2021flow; bengio2023gflownet, genetic algorithms lee2023drug, query-based optimization hoffman2022optimizing, and diffusion models lee2023exploring. Recently, large language model-based methods ye2025drugassist; wang2024efficient have appeared as a promising method for molecule design, when used in combination with other methods like genetic algorithms wang2024efficient.
Genetic algorithms for molecular optimization Genetic algorithms have emerged as a state-of-the-art method for molecule optimization tasks gao2022sample; tripp2023geneticalgorithmsstrongbaselines. They work by mimicking an iterative evolutionary process using operations like mutation and crossover on the molecule representations and allowing favorable candidates to survive to the next generation. Some examples include STONED Nigam_Pollice_Krenn_Gomes_Aspuru-Guzik_2021 which operates on SELFIES representations, GEAM lee2023exploring which operates on molecule fragments, Graph-GA jensen2019graph which applies graph-based mutation/crossover operators for GA, Mol-GA tripp2023geneticalgorithmsstrongbaselines which incorporates quantile uniform sampling to maintain diversity while rewarding the best candidates, MOLLEO wang2024efficient which uses Graph-GA in combination with a large language model, genetic guided GFlowNets kim2024geneticguidedgflownetssampleefficient, and SynNet which incorporates synthesis constraints gao2021amortized. The present work differs from those earlier ones, as it combines a test-time guided generation using a small chemical language model with GA for optimization.
Test-time steering of autoregressive language models Recently, several approaches have been proposed to steer the output of language models to desired outputs without retraining the entire model. Deng. et al deng2023reward proposed Reward-guided Decoding, which uses a reward model to score generations as they are produced and rescales sampling probabilities to favor high-reward tokens. Another approach is Self-disciplined Autoregressive Sampling (SASA) ko2024large, which uses the contextual representations learned from the LLM itself to guide it to generate non-toxic text. Lee et. al lee2024programming, uses conditional activation steering to selectively apply or withhold activation steering using LLM activation patterns. While GP-MoLFormer-Sim also relies on test-time steering of a (chemical) language model, different from the prior works it does not involve training of an external or an internal reward model to be used as guidance during decoding, nor does it require analyzing activation patterns of the decoder. Rather, the proposed method exploits contextual similarity with the target at each step of decoding and updates the logits accordingly.
4 Experiments
4.1 Similarity-guided molecule generation
This task involves generating chemical SMILES similar to a query molecule. We consider five trypsin inhibitors from hartenfeller2012dogs as the targets. The baselines considered are random sampling from a 50k pool of unconditionally generated molecules using GP-MoLFormer, a random search in the reaction template and reactant space until a termination condition is met gottipati2020learningnavigatesyntheticallyaccessible, and a RL-tuned graph isomorphism network (GIN) model that rewards molecules of high similarity to the target guptasimilarity. We report mean Tanimoto similarity ($T_{sim}$), estimated using Morgan fingerprints with a radius of 2, as well as min, max, and mean drug-likeness (QED) over 5 generated sets (one per target) containing the top-k most similar molecules ($k=1,10,10^{2},10^{3}$ and $10^{4}$).
Table 1: Average similarity and QED values of the five trypsin inhibitor targeted generations, generated by various methods.
$T_{sim}$ QED $T_{sim}$ QED
Config top-k mean min mean max Config top-k mean min mean max
$10^{0}$ 1.000 0.225 0.225 0.225 $10^{0}$ 0.694 0.289 0.289 0.289
$10^{1}$ 0.972 0.158 0.238 0.357 $10^{1}$ 0.618 0.12 0.201 0.289
GPMFS $10^{2}$ 0.877 0.075 0.241 0.555 S Modelguptasimilarity $10^{2}$ 0.554 0.049 0.206 0.712
(Ours) $10^{3}$ 0.763 0.037 0.240 0.738 $10^{3}$ 0.499 0.024 0.262 0.923
$10^{4}$ 0.573 0.016 0.213 0.901 $10^{4}$ 0.439 0.013 0.302 0.946
$10^{0}$ 0.438 0.555 0.555 0.555 $10^{0}$ 0.477 0.483 0.483 0.483
Random $10^{1}$ 0.391 0.141 0.491 0.722 Random $10^{1}$ 0.45 0.353 0.448 0.559
Generations $10^{2}$ 0.348 0.048 0.520 0.866 Searchguptasimilarity $10^{4}$ 0.417 0.109 0.418 0.841
$10^{3}$ 0.290 0.019 0.550 0.943 $10^{3}$ 0.377 0.041 0.385 0.841
$10^{4}$ 0.225 0.011 0.571 0.947 $10^{4}$ 0.333 0.022 0.316 0.929
Table 1 reports the mean similarity ($T_{sim}$) of top-k most similar generations obtained using GP-MoLFormer-Sim (GPMFS) and baseline methods. Similarity values of the top-ranked generations show that target similarity-guided decoding using GPMFS performs better than test-time baselines like random sampling and random search. The proposed method also outperforms a graph generative model that is RL-tuned to optimize the target similarity (S model) across all values of k up to $10^{4}$. The reported min, mean, and max QED values show the inverse relation between QED and similarity for these targets, given the mean QED value of these five targets is only 0.234. Nevertheless, 132 molecules are found to have a QED value >0.7>0.7 in the top-10000 most similar molecules generated by GPMFS. Those show a mean similarity of 0.47 and a max similarity of 0.67, showcasing the potential of the proposed method to guide generations on-the-fly toward a target molecule, while yielding useful molecules. A small sample of molecules generated by the GPMFS and their respective targets are visualized in Figure 2 for varying parameter settings (final parameter values can be found in the Appendix Table 17).
4.2 Sample-efficient molecular optimization — PMO benchmark
The open-source benchmark for practical molecular optimization, PMO gao2022sample, has served as an enabler for the transparent and robust evaluation of diverse sets of molecular optimization algorithms. It involves 23 single-objective optimization tasks, that includes property optimization, molecular rediscovery, and structure-based drug design, with a specific focus on the sample efficiency. PMO includes comparing optimization algorithms involving reinforcement learning, Bayesian optimization, generative models, GFlowNets, and genetic algorithms. We compare GP-MoLFormer-Sim+GA (GPMFS+GA) with existing GA-based molecular optimization methods on this benchmark, while focusing on sample efficiency. Following Gao, et al. gao2022sample, we measure the performance by the area under the curve (AUC) of the average property scores of the top-10 molecules versus oracle calls, with the number of maximum oracle calls being 10k. We utilize the task-specific oracles implemented in the Therapeutics Data Commons (TDC) library huang2021therapeuticsdatacommonsmachine. Average and standard deviation of scores obtained from five independent runs starting from different random seeds are reported, unless stated otherwise.
Table 2: Comparison of the guided generation GP-MoLFormer-Sim+GA (“GPMFS+GA”) to selected training-free GA-based baselines. The values for Graph-GA, STONED SELFIES and SynNet are taken from gao2022sample. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Our Rank GPMFS+GA Graph-GA STONED SynNet MOL-GA MOLLEO
(Ours) jensen2019graph SELFIES Nigam_Pollice_Krenn_Gomes_Aspuru-Guzik_2021 Synthesis gao2021amortized tripp2023geneticalgorithmsstrongbaselines (GPT-4)wang2024efficient
albuterol_similarity 4 0.824 (.071) 0.838 (.016) 0.745 (.076) 0.584 (.039) 0.896 (.035) 0.985 (.024)
amlodipine_mpo 3 0.680 (.064) 0.661 (.020) 0.608 (.046) 0.565 (.007) 0.688 (.039) 0.773 (.037)
celecoxib_rediscovery 2 0.716 (.067) 0.630 (.097) 0.382 (.041) 0.441 (.027) 0.567 (.083) 0.864 (.034)
deco_hop 2 0.710 (.058) 0.619 (.004) 0.611 (.008) 0.613 (.009) 0.649 (.025) 0.942 (.013)
DRD2 4 0.956 (.010) 0.964 (.012) 0.913 (.020) 0.969 (.004) 0.936 (.016) 0.968 (.012)
fexofenadine_mpo 3 0.798 (.028) 0.760 (.011) 0.797 (.016) 0.761 (.015) 0.825 (.019) 0.847 (.018)
GSK3 1 0.896 (.035) 0.788 (.070) 0.668 (.049) 0.789 (.032) 0.843 (.039) 0.863 (.047)
isomers_c7h8n2o2 2 0.932 (.011) 0.862 (.065) 0.899 (.011) 0.455 (.031) 0.878 (.026) 0.984 (.008)
isomers_c9h10n2o2pf2cl 3 0.864 (.016) 0.719 (.047) 0.805 (.031) 0.241 (.064) 0.865 (.012) 0.874 (.053)
JNK3 1 0.806 (.087) 0.553 (.136) 0.523 (.092) 0.630 (.034) 0.702 (.123) 0.790 (.027)
median1 2 0.340 (.034) 0.294 (.021) 0.266 (.016) 0.218 (.008) 0.257 (.009) 0.352 (.024)
median2 4 0.255 (.031) 0.273 (.009) 0.245 (.032) 0.235 (.006) 0.301 (.021) 0.275 (.045)
mestranol_similarity 2 0.658 (.118) 0.579 (.022) 0.609 (.101) 0.399 (.021) 0.591 (.053) 0.972 (.009)
osimertinib_mpo 5 0.819 (.004) 0.831 (.005) 0.822 (.012) 0.796 (.003) 0.844 (.015) 0.835 (.024)
perindopril_mpo 2 0.584 (.042) 0.538 (.009) 0.488 (.011) 0.557 (.011) 0.547 (.022) 0.600 (.031)
QED 6 0.940 (.001) 0.940 (.000) 0.941 (.000) 0.941 (.000) 0.941 (.001) 0.948 (.000)
ranolazine_mpo 1 0.812 (.024) 0.728 (.012) 0.765 (.029) 0.741 (.010) 0.804 (.011) 0.769 (.022)
scaffold_hop 2 0.531 (.016) 0.517 (.007) 0.521 (.034) 0.502 (.012) 0.527 (.025) 0.971 (.004)
sitagliptin_mpo 3 0.501 (.081) 0.433 (.075) 0.393 (.083) 0.025 (.014) 0.582 (.040) 0.584 (.067)
thiothixene_rediscovery 3 0.504 (.033) 0.479 (.025) 0.367 (.027) 0.401 (.019) 0.519 (.041) 0.727 (.052)
troglitazone_rediscovery 2 0.437 (.067) 0.390 (.016) 0.320 (.018) 0.283 (.008) 0.427 (.031) 0.562 (.019)
valsartan_smarts 2 0.158 (.317) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.867 (.092)
zaleplon_mpo 3 0.504 (.022) 0.346 (.032) 0.325 (.027) 0.341 (.011) 0.519 (.029) 0.510 (.031)
Average 2.7 0.662 (.221) 0.597 (.233) 0.566 (.245) 0.499 (.259) 0.639 (.236) 0.777 (.200)
Rank by avg. score 2 4 5 6 3 1
Table 2 reports the performance of GPMFS+GA on the 23 tasks from the PMO benchmark. Since the proposed method relies on a combination of the test-time steering of the deep generative model and a modified genetic algorithm, in the main article we show comparison of the proposed method with GA-based baselines specifically designed for molecular design that do not require any training of the generative model. The baselines shown in Table 2 are Graph GA jensen2019graph, STONED Nigam_Pollice_Krenn_Gomes_Aspuru-Guzik_2021, SynNet gao2021amortized, Mol-GA tripp2023geneticalgorithmsstrongbaselines, and MOLLEO wang2024efficient. We report the rank per task based on the top-10 AUC score obtained with a maximum of 10k oracle calls for each method. Average rank and average score over all tasks are also reported in Table 2. For comparison with additional baselines and more analyses on the optimization runs, see Appendix. Results show that the proposed method scores second among GA baselines, while MOLLEO that uses Graph-GA with GPT-4 comes first. On three tasks, namely GSK3, JNK3, and ranolazine_mpo, GPMFS+GA outperforms all baselines, while on another 9 tasks it ranks second. Given the computational and actual dollar cost associated with calling GPT-4, GPMFS+GA appears as a more cost-effective alternative. To gauge synthesizability, average Synthetic Accessibility scores of the top 100 generated molecules in each task are reported in Appendix Table 16.
We also compare GPMFS+GA with other MOLLEO variants that use a smaller domain-aware language model — namely, a BioT5 model and a MoleculeSTM model (see Appendix Table 15). GPMFS+GA performs better than MOLLEO (MoleculeSTM), while losing against the BioT5 variant.
Table 3: Comparison of GP-MoLFormer-Sim+GA with MOLLEO in the black-box oracle setting. Molecule name in prompt is redacted for MOLLEO in that setting. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
black-box
Task GPMFS+GA MOLLEO MOLLEO
(ours) (GPT-4.1-mini, redacted) (GPT-4.1-mini)
thiothixene_rediscovery 0.504 (.033) 0.462 (.031) 0.692 (.013)
mestranol_similarity 0.658 (.118) 0.644 (.065) 0.983 (.001)
4.3 Comparison with MOLLEO in the black-box oracle setting
MOLLEO includes natural language prompting of the LLM to generate proposals based on the GA operations — crossover and mutation. While doing so, the prompt includes information about the task and the oracle function. For example, for the thiothixene rediscovery task, the prompt used in reference wang2024efficient includes the following (emphasis ours): “OBJECTIVE: has a higher thiothixene rediscovery score. TASK: thiothixene rediscovery scores. OBJECTIVE_DEFINITION: The thiothixene rediscovery score measures a molecule’s Tanimoto similarity with thiothixene’s SMILES to check whether it could be rediscovered.” This contextual information present in the prompt weakens the black-box nature of the oracle used in the PMO benchmark as the GPT-4 model has memorized the thiothixene SMILES, which is otherwise never disclosed to the other baselines. Therefore, to enable a fair comparison, we revise the prompt such that the name of the target molecule is redacted from the prompt. Results are reported in Table 3 for two exemplar tasks, namely thiothixene rediscovery and mestranol similarity. On both tasks, MOLLEO’s performance (using a gpt-4.1-mini model) drops by $\sim$33% and becomes worse compared to GPMFS+GA when the molecule name is redacted. This result implies that MOLLEO’s performance depends on the LLM’s utilization of the task-relevant contextual information for proposing offspring. In contrast, our proposed method only uses the score from the (black-box) oracle, consistent with the setting of the PMO benchmark, to produce candidates during optimization and outperforms MOLLEO in that mode.
We also run an experiment where GPMFS+GA has access to the target SMILES information used in the oracle function (when applicable) and utilizes that to create the initial pool of candidates for optimization. Table 4 shows the performance gain achieved by the proposed method in that mode, again underscoring the inflationary effect of breaking the black-box nature of the oracle on the performance.
Table 4: Top-10 AUC metrics for GP-MoLFormer-Sim+GA (“GPMFS+GA”) under access to the oracle’s target SMILES in comparison to MOLLEO models. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Our GPMFS+GA MOLLEO MOLLEO MOLLEOwang2024efficient
Rank (Ours) (MolSTM) (BioT5) (GPT-4)
albuterol 1 0.994 (.005) 0.929 (.005) 0.968 (.003) 0.985 (.024)
amlodipine 3 0.719 (.061) 0.674 (.018) 0.776 (.038) 0.773 (.037)
celecoxib 1 0.885 (.010) 0.594 (.105) 0.508 (.017) 0.864 (.034)
fexofenadine 2 0.812 (.046) 0.789 (.016) 0.773 (.017) 0.847 (.018)
median1 1 0.405 (.001) 0.298 (.019) 0.338 (.033) 0.352 (.024)
median2 1 0.393 (.004) 0.251 (.031) 0.259 (.019) 0.275 (.045)
mestranol 1 0.993 (.011) 0.596 (.018) 0.717 (.104) 0.972 (.009)
osimertinib 3 0.818 (.002) 0.823 (.007) 0.817 (.016) 0.835 (.024)
perindopril 3 0.582 (.018) 0.554 (.037) 0.738 (.016) 0.600 (.031)
ranolazine 1 0.825 (.016) 0.725 (.040) 0.749 (.012) 0.769 (.022)
sitagliptin 4 0.435 (.054) 0.548 (.065) 0.506 (.100) 0.584 (.067)
thiothixene 1 0.862 (.020) 0.508 (.035) 0.696 (.081) 0.727 (.052)
troglitazone 1 0.905 (.002) 0.381 (.025) 0.390 (.044) 0.562 (.019)
zaleplon 1 0.684 (.024) 0.475 (.018) 0.465 (.026) 0.510 (.031)
Average 1.7 0.737 (.201) 0.582 (.189) 0.621 (.201) 0.690 (.209)
Rank by avg. score 1 4 3 2
4.4 Ablation experiments
Algorithm 2 in Section 2.3 provides for several optional steps, namely: (1) using random Fourier features (RFF) to approximate a kernel distance in the GP-MoLFormer embedding space, (2) using a genetic graph-based crossover (XO) between parent molecules (drawn from the set of best guides), and (3) adding a set of diverse guides (DIV) to enhance exploration. Results reported in Tables 2 and 4 were obtained employing all of these variants active. To tease apart individual effects of these options, we also ran a series of ablation experiments over the 23 PMO tasks. Table 5 summarizes the relative performance in multiple configurations, starting with guided generation (GG) with none of the three options to GG with the full set active. For each combination we report the average rank over all tasks as well as the average AUC metric. We observe each option adding a benefit, with the exception of the RFF, however, only when being added alone. The best configuration is GG+RFF768+XO+DIV which is used throughout our PMO experiments, unless otherwise stated. Detailed, per-task ablation results are given in the Appendix (Table 7) along with further hyperparameter details in Table 17.
Table 5: Overall improvement due to specific method combinations in terms of average 1-based rank and top-10 AUC metric in GP-MoLFormer-Sim. Includes adding “XO” as crossover, “RFF768” as 768-dimensional Random Fourier Features, and “DIV” as diversity augmentation to the vanilla Guided Generation ($G$).
Config Avg. rank $\downarrow$ Avg. score $\uparrow$
$G$ 5.0 0.603
+XO 3.4 0.672
+RFF768 5.3 0.597
+XO+DIV 3.0 0.678
+RFF768+XO 2.5 0.682
+RFF768+XO+DIV 1.8 0.690
5 Limitations and Broader Impact
Given the need for discovering new and useful artifacts for various discovery applications, the proposed method can have broader impact beyond chemistry and biology. There remain open questions and limitations, however. For example, although the proposed framework is model and domain-agnostic, we have only experimented here with a specific chemical language model decoder and on specific molecular optimization tasks. It also remains an open question to what extent the method can benefit from including “negative” targets while subjected to an optimization task. Extending the method to multi-objective optimization (optimizing linear combination of multiple objectives with different importances) can also be a potential future research direction.
6 Conclusions
In this work, we present GP-MoLFormer-Sim, a test-time framework for sequentially revising the generated output of a small chemical language model to maintain the contextual closeness between its generation and a given set of targets. Furthermore, we integrate the proposed method into a genetic algorithm as an effective proposer of mutations to produce high-quality offspring (GP-MoLFormer-Sim+GA). Our method is validated on a variety of molecular optimization tasks. Evaluating this framework with a black-box oracle reveals performance improvement compared to a baseline that leverages large language model like GPT-4, demonstrating the effective trade-off between performance and computational efficiency of the proposed method. We believe the proposed guided generation method represents a versatile and valuable addition to the modeling toolbox in molecular optimization and beyond.
Appendix A Theoretical Justification for Algorithm 1
Let $V$ be the vocabulary size. Given previously generated tokens represented as context $c$, let $\pi^{\text{target}}_{\mathsf{KDE}}(\cdot|c)$ be the Kernel Density Estimator (KDE) of the target molecules defined in the embedding space of GP-MoLFormer noted below as $\pi_{\mathrm{ref}}$. Our algorithm can be seen at each token generation as the closed form solution of the following optimization problem:
$\mathsf{P:}\max_{p\in\Delta_{V}}$ $(1-\alpha)\underbrace{\sum_{i=1}^{V}p_{i}\pi^{\text{target}}_{\mathrm{KDE}}(x_{i}|c)}_{\text{Similarity to target }\textsf{KDE}}-\underbrace{\mathsf{KL}(p||\pi_{\mathrm{ref}}^{\alpha}(\cdot|c))}_{\mathsf{KL}\textsf{ to {GP-MoLFormer}}}$
$\mathsf{s.t.}~\Delta_{V}=\left\{p\in[0,1]^{V}\Big|\sum_{i=1}^{V}p_{i}=1\right\},$
where for $p\in\Delta_{V}$ and $q\in\mathbb{R}_{+}^{V},$ we define the generalized KL divergence [peyré2020computationaloptimaltransport] as:
$\mathrm{KL}(p\,\|\,q)=\sum_{i=1}^{V}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)+\sum_{i=1}^{V}q_{i}-1.$
The optimization problems $\mathsf{P}$ balances between closeness to GP-MoLFormer distribution smoothed by $\alpha$ ($\pi_{\mathrm{ref}}^{\alpha}$) and the similarity to the target molecules kernel density estimator ($\pi^{\text{target}}_{\mathrm{KDE}}$). We use the notion of generalized divergence since $\pi_{\mathrm{ref}}^{\alpha}$ is a smoothed distribution that is no longer normalized. Indeed the optimal solution of this problem [ko2025large] is proportional to $\pi_{\mathrm{ref}}^{\alpha}\exp((1-\alpha)\pi^{\text{target}}_{\mathrm{KDE}})=\exp((1-\alpha)\pi^{\text{target}}_{\mathrm{KDE}}+\alpha\log\pi_{\mathrm{ref}})$. This justifies the computations on lines 23–24 in Algorithm 1, where the new guided probability is obtained using a softmax on the interpolation between the KDE estimator of the target molecules and the logits of GP-MoLFormer.
Appendix B Extended Results — PMO Tasks
B.1 Optimization trajectory — an example
Figure 3 visualizes the GP-MoLFormer-Sim+GA process in the GP-MoLFormer embedding space (projected onto a 2D t-SNE plane). Starting with a random subset of 100 ZINC molecules (blue dots in the center), the optimization picks up on a handful highest-scoring candidates (molecular structure of one of these is depicted at the top right) and through the guided generation comes up with offspring displayed as dot in orange color (marked as generation 2). The GA process continues a milestones are numbered in Figure 3 by their generation number and shown along with representative chemical structures. The process completes on generation 252 upon exhausting the oracle budget of 10000, reaching a final oracle value of 0.897. Also shown is the actual oracle target, the compound albuterol. Note that only a single guide trajectory is shown in this figure for sake of simplicity. In our albuterol runs reported in Table 2 and Figure 4, the actual target was hit exactly.
Figure 3: Albuterol optimization trajectory of a single guide visualized in GP-MoLFormer space via a 2D t-SNE projection.
B.2 Redacted MOLLEO Prompts
Table 6 lists the modified prompts used for gauging performance gain/loss due to knowledge of the oracle target compound (Section 4.3).
Table 6: Prompts used for redacted MOLLEO. Words in italics are modified from the original prompts.
Task Description Objective
thiothixene_rediscovery I have two molecules and their rediscovery score measures a molecule’s Tanimoto similarity with a particular SMILES to check whether it could be rediscovered. Please propose a new molecule that has a higher rediscovery score.
mestranol_similarity I have two molecules and their target similarity scores. The target similarity score measures a molecule’s Tanimoto similarity with a particular target molecule. Please propose a new molecule that has a higher target similarity score.
B.3 Configuration ablation
Table 7 gives efficacy details regarding individual features of the GPMoLFormer-Sim+GA procedure, including adding (1) 768-dimensional random Fourier features (“RFF768”), (2) Crossover (“XO”), and Diversity guides (“DIV”). Overall, the full combination +RFF768+XO+DIV gives best results, as can also be seen in the aggregate score in the last row of the table.
Table 7: Top-10 AUC PMO metrics obtained using various configurations (mean ($\pm$ standard deviation) over 5 runs for each).
Task Guided Gen. +XO +RFF768 +XO+DIV +RFF768+XO +RFF768+XO+DIV
albuterol 0.714 (.044) 0.803 (.045) 0.767 (.067) 0.847 (.050) 0.814 (.039) 0.824 (.071)
amlodipine 0.587 (.034) 0.699 (.036) 0.584 (.006) 0.625 (.022) 0.715 (.067) 0.680 (.064)
amlomestranol 0.542 (.059) 0.579 (.072) 0.542 (.034) 0.602 (.049) 0.603 (.054) 0.658 (.118)
celecoxib 0.560 (.114) 0.627 (.052) 0.577 (.141) 0.667 (.079) 0.758 (.070) 0.716 (.067)
deco_hop 0.748 (.113) 0.701 (.086) 0.668 (.088) 0.649 (.011) 0.629 (.009) 0.710 (.058)
DRD2 0.929 (.037) 0.955 (.021) 0.954 (.026) 0.958 (.018) 0.952 (.015) 0.956 (.010)
fexofenadine 0.756 (.020) 0.791 (.012) 0.801 (.031) 0.802 (.020) 0.840 (.030) 0.798 (.028)
GSK3 0.818 (.061) 0.833 (.073) 0.781 (.070) 0.862 (.081) 0.871 (.066) 0.896 (.035)
isomers_c7 0.916 (.028) 0.921 (.026) 0.907 (.035) 0.912 (.011) 0.933 (.006) 0.932 (.011)
isomers_c9 0.819 (.025) 0.842 (.027) 0.831 (.034) 0.847 (.023) 0.861 (.015) 0.864 (.016)
JNK3 0.654 (.037) 0.773 (.085) 0.609 (.078) 0.791 (.038) 0.801 (.092) 0.806 (.087)
median1 0.277 (.011) 0.352 (.035) 0.320 (.032) 0.358 (.035) 0.369 (.022) 0.340 (.034)
median2 0.229 (.012) 0.218 (.012) 0.204 (.013) 0.252 (.027) 0.234 (.028) 0.255 (.031)
osimertimib 0.806 (.018) 0.813 (.011) 0.807 (.016) 0.812 (.007) 0.815 (.009) 0.819 (.004)
perindopril 0.531 (.027) 0.580 (.037) 0.528 (.033) 0.551 (.014) 0.572 (.036) 0.584 (.042)
QED 0.939 (.002) 0.940 (.001) 0.939 (.001) 0.940 (.001) 0.941 (.000) 0.940 (.001)
ranolazine 0.782 (.012) 0.805 (.009) 0.799 (.021) 0.788 (.016) 0.815 (.007) 0.812 (.024)
scaffold_hop 0.519 (.019) 0.543 (.018) 0.511 (.016) 0.527 (.004) 0.539 (.012) 0.531 (.016)
sitagliptin 0.364 (.073) 0.451 (.010) 0.384 (.041) 0.480 (.032) 0.445 (.050) 0.501 (.081)
thiothixene 0.429 (.008) 0.526 (.037) 0.432 (.008) 0.525 (.012) 0.461 (.011) 0.504 (.033)
troglitazone 0.346 (.059) 0.385 (.090) 0.329 (.039) 0.408 (.098) 0.420 (.051) 0.437 (.067)
valsartan_smarts 0.100 (.199) 0.040 (.080) 0.000 (.000) 0.118 (.180) 0.073 (.087) 0.158 (.317)
zaleplon 0.496 (.029) 0.493 (.007) 0.469 (.014) 0.497 (.017) 0.487 (.017) 0.504 (.022)
Avg. Score 0.603 0.638 0.597 0.644 0.650 0.662
B.3.1 Optimization curves
Figure 4 shows how the GP-MoLFormer-Sim+GA optimization progresses as a function of number of Oracle calls in the 23 PMO tasks (grouped by type). We observe a variety of patterns: while the similarity, rediscovery and MPO task curves look fairly similar in range and progression, the PO tasks (specifically, QED and DRD2) tend to quickly ramp up with small variance across the 5 runs. On the other hand, the median tasks seem to be significantly more challenging and the process seems to quickly plateau after a few hundred oracle calls. This reflects the nature of the median task and is not unexpected. Finally, the SBO tasks show the largest variability with the special case of valsartan - a task on which most baselines fail to obtain any hits (usually remain flat at 0.0). We believe that our valsartan task result shown in Figure 4 highlights the exploratory power of the crossover and diversity enhanced guided generation.
Figure 4: Optimization curves for the PMO tasks by group.
B.3.2 Extended baseline comparison
We list detailed comparison of the GP-MoLFormer-Sim GA top-10 AUC metrics with baselines published previously.
Table 8 shows a comparison of our GPMFS-GA to [tripp2023geneticalgorithmsstrongbaselines]. Tables 9 through 12 show results reported in [gao2022sample]. Tables 13-14 compare results from [kim2024geneticguidedgflownetssampleefficient], and Table 15 compares results published in [wang2024efficient].
Table 8: Comparison of GP-MoLFormer-Sim+GA to main results in [tripp2023geneticalgorithmsstrongbaselines]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Our Rank Our AUC REINVENT Graph GA MOL_GA
albuterol 4 0.824 (.071) 0.882 (.006) 0.838 (.016) 0.896 (.035)
amlodipine 2 0.680 (.064) 0.635 (.035) 0.661 (.020) 0.688 (.039)
celecoxib 1 0.716 (.067) 0.713 (.067) 0.630 (.097) 0.567 (.083)
deco_hop 1 0.710 (.058) 0.666 (.044) 0.619 (.004) 0.649 (.025)
DRD2 2 0.956 (.010) 0.945 (.007) 0.964 (.012) 0.936 (.016)
fexofenadine 2 0.798 (.028) 0.784 (.006) 0.760 (.011) 0.825 (.019)
GSK3 1 0.896 (.035) 0.865 (.043) 0.788 (.070) 0.843 (.039)
isomers_c7 1 0.932 (.011) 0.852 (.036) 0.862 (.065) 0.878 (.026)
isomers_c9 2 0.864 (.016) 0.642 (.054) 0.719 (.047) 0.865 (.012)
JNK3 1 0.806 (.087) 0.783 (.023) 0.553 (.136) 0.702 (.123)
median1 2 0.340 (.034) 0.356 (.009) 0.294 (.021) 0.257 (.009)
median2 4 0.255 (.031) 0.276 (.008) 0.273 (.009) 0.301 (.021)
mestranol 1 0.658 (.118) 0.618 (.048) 0.579 (.022) 0.591 (.053)
osimertinib 4 0.819 (.004) 0.837 (.009) 0.831 (.005) 0.844 (.015)
perindopril 1 0.584 (.042) 0.537 (.016) 0.538 (.009) 0.547 (.022)
QED 3 0.940 (.001) 0.941 (.000) 0.940 (.000) 0.941 (.001)
ranolazine 1 0.812 (.024) 0.760 (.009) 0.728 (.012) 0.804 (.011)
scaffold_hop 2 0.531 (.016) 0.560 (.019) 0.517 (.007) 0.527 (.025)
sitagliptin 2 0.501 (.081) 0.021 (.003) 0.433 (.075) 0.582 (.040)
thiothixene 3 0.504 (.033) 0.534 (.013) 0.479 (.025) 0.519 (.041)
troglitazone 2 0.437 (.067) 0.441 (.032) 0.390 (.016) 0.427 (.031)
valsartan_smarts 2 0.158 (.317) 0.178 (.358) 0.000 (.000) 0.000 (.000)
zaleplon 2 0.504 (.022) 0.358 (.062) 0.346 (.032) 0.519 (.029)
Average 2.0 (1.0) 0.662 (.221) 0.617 (.245) 0.597 (.233) 0.639 (.236)
Rank by avg. score 1 3 4 2
Table 9: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 1 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Rank Our REINVENT Graph GA REINVENT GP BO STONED LSTM HC SMILES
Rank AUC SMILES Fragments SELFIES Fragments SELFIES SMILES GA
albuterol 5 0.824 (.071) 0.882 (.006) 0.838 (.016) 0.826 (.030) 0.898 (.014) 0.745 (.076) 0.719 (.018) 0.661 (.066)
amlodipine 1 0.680 (.064) 0.635 (.035) 0.661 (.020) 0.607 (.014) 0.583 (.044) 0.608 (.046) 0.593 (.016) 0.549 (.009)
celecoxib 2 0.716 (.067) 0.713 (.067) 0.630 (.097) 0.573 (.043) 0.723 (.053) 0.382 (.041) 0.539 (.018) 0.344 (.027)
deco_hop 4 0.710 (.058) 0.666 (.044) 0.619 (.004) 0.631 (.012) 0.629 (.018) 0.611 (.008) 0.826 (.017) 0.611 (.006)
DRD2 3 0.956 (.010) 0.945 (.007) 0.964 (.012) 0.943 (.005) 0.923 (.017) 0.913 (.020) 0.919 (.015) 0.908 (.019)
fexofenadine 1 0.798 (.028) 0.784 (.006) 0.760 (.011) 0.741 (.002) 0.722 (.005) 0.797 (.016) 0.725 (.003) 0.721 (.015)
GSK3 1 0.896 (.035) 0.865 (.043) 0.788 (.070) 0.780 (.037) 0.851 (.041) 0.668 (.049) 0.839 (.015) 0.629 (.044)
isomers_c7 1 0.932 (.011) 0.852 (.036) 0.862 (.065) 0.849 (.034) 0.680 (.117) 0.899 (.011) 0.485 (.045) 0.913 (.021)
isomers_c9 1 0.864 (.016) 0.642 (.054) 0.719 (.047) 0.733 (.029) 0.469 (.180) 0.805 (.031) 0.342 (.027) 0.860 (.065)
JNK3 1 0.806 (.087) 0.783 (.023) 0.553 (.136) 0.631 (.064) 0.564 (.155) 0.523 (.092) 0.661 (.039) 0.316 (.022)
median1 3 0.340 (.034) 0.356 (.009) 0.294 (.021) 0.355 (.011) 0.301 (.014) 0.266 (.016) 0.255 (.010) 0.192 (.012)
median2 4 0.255 (.031) 0.276 (.008) 0.273 (.009) 0.255 (.005) 0.297 (.009) 0.245 (.032) 0.248 (.008) 0.198 (.005)
mestranol 1 0.658 (.118) 0.618 (.048) 0.579 (.022) 0.620 (.029) 0.627 (.089) 0.609 (.101) 0.526 (.032) 0.469 (.029)
osimertinib 5 0.819 (.004) 0.837 (.009) 0.831 (.005) 0.820 (.003) 0.787 (.006) 0.822 (.012) 0.796 (.002) 0.817 (.011)
perindopril 1 0.584 (.042) 0.537 (.016) 0.538 (.009) 0.517 (.021) 0.493 (.011) 0.488 (.011) 0.489 (.007) 0.447 (.013)
QED 4 0.940 (.001) 0.941 (.000) 0.940 (.000) 0.940 (.000) 0.937 (.000) 0.941 (.000) 0.939 (.000) 0.940 (.000)
ranolazine 1 0.812 (.024) 0.760 (.009) 0.728 (.012) 0.748 (.018) 0.735 (.013) 0.765 (.029) 0.714 (.008) 0.699 (.026)
scaffold_hop 4 0.531 (.016) 0.560 (.019) 0.517 (.007) 0.525 (.013) 0.548 (.019) 0.521 (.034) 0.533 (.012) 0.494 (.011)
sitagliptin 1 0.501 (.081) 0.021 (.003) 0.433 (.075) 0.194 (.121) 0.186 (.055) 0.393 (.083) 0.066 (.019) 0.363 (.057)
thiothixene 3 0.504 (.033) 0.534 (.013) 0.479 (.025) 0.495 (.040) 0.559 (.027) 0.367 (.027) 0.438 (.008) 0.315 (.017)
troglitazone 2 0.437 (.067) 0.441 (.032) 0.390 (.016) 0.348 (.012) 0.410 (.015) 0.320 (.018) 0.354 (.016) 0.263 (.024)
valsartan 2 0.158 (.317) 0.179 (.358) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000)
zaleplon 1 0.504 (.022) 0.358 (.062) 0.346 (.032) 0.333 (.026) 0.221 (.072) 0.325 (.027) 0.206 (.006) 0.334 (.041)
Average 2.3 (1.4) 0.662 (.221) 0.617 (.245) 0.597 (.233) 0.585 (.242) 0.571 (.244) 0.566 (.245) 0.531 (.258) 0.524 (.258)
Rank by
avg. score 1 2 3 4 5 6 7 8
Table 10: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 2 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task SynNet Synthesis DoG-Gen Synthesis DST Fragments MARS MIMOSA MolPal LSTM HC SELFIES
albuterol 0.584 (.039) 0.676 (.013) 0.619 (.020) 0.597 (.124) 0.618 (.017) 0.609 (.002) 0.664 (.030)
amlodipine 0.565 (.007) 0.536 (.003) 0.516 (.007) 0.504 (.016) 0.543 (.003) 0.582 (.008) 0.532 (.004)
celecoxib 0.441 (.027) 0.464 (.009) 0.380 (.006) 0.379 (.060) 0.393 (.010) 0.415 (.001) 0.385 (.008)
deco_hop 0.613 (.009) 0.800 (.007) 0.608 (.008) 0.589 (.003) 0.619 (.003) 0.643 (.005) 0.590 (.001)
DRD2 0.969 (.004) 0.948 (.001) 0.820 (.014) 0.891 (.020) 0.799 (.017) 0.783 (.009) 0.729 (.034)
fexofenadine 0.761 (.015) 0.695 (.003) 0.725 (.005) 0.711 (.006) 0.706 (.011) 0.685 (.000) 0.693 (.004)
GSK3 0.789 (.032) 0.831 (.021) 0.671 (.032) 0.552 (.037) 0.554 (.042) 0.555 (.011) 0.423 (.018)
isomers_c7 0.455 (.031) 0.465 (.018) 0.548 (.069) 0.728 (.027) 0.564 (.046) 0.484 (.006) 0.587 (.031)
isomers_c9 0.241 (.064) 0.199 (.016) 0.458 (.063) 0.581 (.013) 0.303 (.046) 0.164 (.003) 0.352 (.019)
JNK3 0.630 (.034) 0.595 (.023) 0.556 (.057) 0.489 (.095) 0.360 (.063) 0.339 (.009) 0.207 (.013)
median1 0.218 (.008) 0.217 (.001) 0.232 (.009) 0.207 (.011) 0.243 (.005) 0.249 (.001) 0.239 (.009)
median2 0.235 (.006) 0.212 (.000) 0.185 (.020) 0.181 (.011) 0.214 (.002) 0.230 (.000) 0.205 (.005)
mestranol 0.399 (.021) 0.437 (.007) 0.450 (.027) 0.388 (.026) 0.438 (.015) 0.564 (.004) 0.446 (.009)
osimertinib 0.796 (.003) 0.774 (.002) 0.785 (.004) 0.777 (.006) 0.788 (.014) 0.779 (.000) 0.780 (.005)
perindopril 0.557 (.011) 0.474 (.002) 0.462 (.008) 0.462 (.006) 0.490 (.011) 0.467 (.002) 0.448 (.006)
QED 0.941 (.000) 0.934 (.000) 0.938 (.000) 0.930 (.003) 0.939 (.000) 0.940 (.000) 0.938 (.000)
ranolazine 0.741 (.010) 0.711 (.006) 0.632 (.054) 0.740 (.010) 0.640 (.015) 0.457 (.005) 0.614 (.010)
scaffold_hop 0.502 (.012) 0.515 (.005) 0.497 (.004) 0.469 (.004) 0.507 (.015) 0.494 (.000) 0.472 (.002)
sitagliptin 0.025 (.014) 0.048 (.008) 0.075 (.032) 0.016 (.003) 0.102 (.023) 0.043 (.001) 0.116 (.012)
thiothixene 0.401 (.019) 0.375 (.004) 0.366 (.006) 0.344 (.022) 0.347 (.018) 0.339 (.001) 0.339 (.009)
troglitazone 0.283 (.008) 0.416 (.019) 0.279 (.019) 0.256 (.016) 0.299 (.009) 0.268 (.000) 0.257 (.002)
valsartan_smarts 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000)
zaleplon 0.341 (.011) 0.123 (.016) 0.176 (.045) 0.187 (.046) 0.172 (.036) 0.168 (.003) 0.218 (.020)
Average 0.499 (.259) 0.498 (.269) 0.477 (.236) 0.477 (.253) 0.463 (.232) 0.446 (.238) 0.445 (.228)
Rank by avg. score 9 10 12 11 13 14 15
Table 11: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 3 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task DoG-AE GFlowNet Fragments GA+D SELFIES VAE BO SELFIES Screening VAE BO SMILES Pasithea SELFIES
albuterol 0.533 (.034) 0.447 (.012) 0.495 (.025) 0.494 (.012) 0.483 (.006) 0.489 (.007) 0.447 (.007)
amlodipine 0.507 (.005) 0.444 (.004) 0.400 (.032) 0.516 (.005) 0.535 (.001) 0.533 (.009) 0.504 (.003)
celecoxib 0.355 (.012) 0.327 (.004) 0.223 (.025) 0.326 (.007) 0.351 (.005) 0.354 (.002) 0.312 (.007)
deco_hop 0.765 (.055) 0.583 (.002) 0.550 (.005) 0.579 (.001) 0.590 (.001) 0.589 (.001) 0.579 (.001)
DRD2 0.943 (.009) 0.590 (.070) 0.382 (.205) 0.569 (.039) 0.545 (.015) 0.555 (.043) 0.255 (.040)
fexofenadine 0.679 (.017) 0.693 (.006) 0.587 (.007) 0.670 (.004) 0.666 (.004) 0.671 (.003) 0.660 (.015)
GSK3 0.601 (.091) 0.651 (.026) 0.342 (.019) 0.350 (.034) 0.438 (.034) 0.386 (.006) 0.281 (.038)
isomers_c7 0.239 (.077) 0.366 (.043) 0.854 (.015) 0.325 (.028) 0.168 (.034) 0.161 (.017) 0.673 (.030)
isomers_c9 0.049 (.015) 0.110 (.031) 0.657 (.020) 0.200 (.030) 0.106 (.021) 0.084 (.009) 0.345 (.145)
JNK3 0.469 (.138) 0.440 (.022) 0.219 (.021) 0.208 (.022) 0.238 (.024) 0.241 (.026) 0.154 (.018)
median1 0.171 (.009) 0.202 (.004) 0.180 (.009) 0.201 (.003) 0.205 (.005) 0.202 (.006) 0.178 (.009)
median2 0.182 (.006) 0.180 (.000) 0.121 (.005) 0.185 (.001) 0.200 (.004) 0.195 (.001) 0.179 (.004)
mestranol 0.370 (.014) 0.322 (.007) 0.371 (.016) 0.386 (.009) 0.409 (.019) 0.399 (.005) 0.361 (.016)
osimertinib 0.750 (.012) 0.784 (.001) 0.672 (.027) 0.765 (.002) 0.764 (.001) 0.771 (.002) 0.749 (.007)
perindopril 0.432 (.013) 0.430 (.010) 0.172 (.088) 0.429 (.003) 0.445 (.004) 0.442 (.004) 0.421 (.008)
QED 0.926 (.003) 0.921 (.004) 0.860 (.014) 0.936 (.001) 0.938 (.000) 0.938 (.000) 0.931 (.002)
ranolazine 0.689 (.015) 0.652 (.002) 0.555 (.015) 0.452 (.025) 0.411 (.010) 0.457 (.012) 0.347 (.012)
scaffold_hop 0.489 (.010) 0.463 (.002) 0.413 (.009) 0.455 (.004) 0.471 (.002) 0.470 (.003) 0.456 (.003)
sitagliptin 0.009 (.005) 0.008 (.003) 0.281 (.022) 0.084 (.015) 0.022 (.003) 0.023 (.004) 0.088 (.013)
thiothixene 0.314 (.015) 0.285 (.012) 0.223 (.029) 0.297 (.004) 0.317 (.003) 0.317 (.007) 0.288 (.006)
troglitazone 0.259 (.016) 0.188 (.001) 0.152 (.013) 0.243 (.004) 0.249 (.003) 0.257 (.003) 0.240 (.002)
valsartan_smarts 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.002 (.003) 0.000 (.000) 0.002 (.004) 0.006 (.012)
zaleplon 0.049 (.027) 0.035 (.030) 0.244 (.015) 0.206 (.015) 0.072 (.014) 0.039 (.012) 0.091 (.013)
Average 0.425 (.278) 0.397 (.247) 0.389 (.227) 0.386 (.217) 0.375 (.234) 0.373 (.238) 0.372 (.226)
Rank by avg. score 16 17 18 19 20 21 22
Table 12: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 4 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task GFlowNet-AL Fragments JT-VAE BO Fragments Graph MCTS Atoms MolDQN Atoms
albuterol 0.390 (.008) 0.485 (.029) 0.580 (.023) 0.320 (.015)
amlodipine 0.428 (.002) 0.519 (.009) 0.447 (.008) 0.311 (.008)
celecoxib 0.257 (.003) 0.299 (.009) 0.264 (.013) 0.099 (.005)
deco_hop 0.583 (.001) 0.585 (.002) 0.554 (.002) 0.546 (.001)
DRD2 0.468 (.046) 0.506 (.136) 0.300 (.050) 0.025 (.001)
fexofenadine 0.688 (.002) 0.667 (.010) 0.574 (.009) 0.478 (.012)
GSK3 0.588 (.015) 0.350 (.051) 0.281 (.022) 0.241 (.008)
isomers_c7 0.241 (.055) 0.103 (.016) 0.530 (.035) 0.431 (.035)
isomers_c9 0.064 (.012) 0.090 (.035) 0.454 (.067) 0.342 (.026)
JNK3 0.362 (.021) 0.222 (.009) 0.110 (.019) 0.111 (.008)
median1 0.190 (.002) 0.179 (.003) 0.195 (.005) 0.122 (.007)
median2 0.173 (.001) 0.180 (.003) 0.132 (.002) 0.088 (.003)
mestranol 0.295 (.004) 0.356 (.013) 0.281 (.008) 0.188 (.007)
osimertinib 0.787 (.003) 0.775 (.004) 0.700 (.004) 0.674 (.006)
perindopril 0.421 (.002) 0.430 (.009) 0.277 (.013) 0.213 (.043)
QED 0.902 (.005) 0.934 (.002) 0.892 (.006) 0.731 (.018)
ranolazine 0.632 (.007) 0.508 (.055) 0.239 (.027) 0.051 (.020)
scaffold_hop 0.460 (.002) 0.470 (.005) 0.412 (.003) 0.405 (.004)
sitagliptin 0.006 (.001) 0.046 (.027) 0.056 (.012) 0.003 (.002)
thiothixene 0.266 (.005) 0.282 (.008) 0.231 (.004) 0.099 (.007)
troglitazone 0.186 (.003) 0.237 (.005) 0.224 (.009) 0.122 (.004)
valsartan_smarts 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000)
zaleplon 0.010 (.001) 0.125 (.038) 0.058 (.019) 0.010 (.005)
Average 0.365 (.246) 0.363 (.236) 0.339 (.219) 0.244 (.211)
Rank by avg. score 23 24 25 26
Table 13: Comparison of GP-MoLFormer-Sim+GA to main results in [kim2024geneticguidedgflownetssampleefficient] - Page 1 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Our Rank Our AUC Genetic GFN Mol GA SMILES REINVENT GEGL
albuterol 6 0.824 (.071) 0.949 (.010) 0.928 (.015) 0.881 (.016) 0.842 (.019)
amlodipine 3 0.680 (.064) 0.761 (.019) 0.740 (.055) 0.644 (.019) 0.626 (.018)
celecoxib 4 0.716 (.067) 0.802 (.029) 0.629 (.062) 0.717 (.027) 0.699 (.041)
deco_hop 2 0.710 (.058) 0.733 (.109) 0.656 (.013) 0.662 (.044) 0.656 (.039)
DRD2 3 0.956 (.010) 0.974 (.006) 0.950 (.004) 0.957 (.007) 0.898 (.015)
fexofenadine 3 0.798 (.028) 0.856 (.039) 0.835 (.012) 0.781 (.013) 0.769 (.009)
GSK3 1 0.896 (.035) 0.881 (.042) 0.894 (.025) 0.885 (.031) 0.816 (.027)
isomers_c7 3 0.932 (.011) 0.969 (.003) 0.926 (.014) 0.942 (.012) 0.930 (.011)
isomers_c9 3 0.864 (.016) 0.897 (.007) 0.894 (.005) 0.838 (.030) 0.808 (.007)
JNK3 2 0.806 (.087) 0.764 (.069) 0.835 (.040) 0.782 (.029) 0.580 (.086)
median1 3 0.340 (.034) 0.379 (.010) 0.329 (.006) 0.363 (.011) 0.338 (.016)
median2 6 0.255 (.031) 0.294 (.007) 0.284 (.035) 0.281 (.002) 0.274 (.007)
mestranol 3 0.658 (.118) 0.708 (.057) 0.762 (.048) 0.634 (.042) 0.599 (.035)
osimertinib 5 0.819 (.004) 0.860 (.008) 0.853 (.005) 0.834 (.010) 0.832 (.005)
perindopril 3 0.584 (.042) 0.595 (.014) 0.610 (.038) 0.535 (.015) 0.537 (.015)
QED 5 0.940 (.001) 0.942 (.000) 0.941 (.001) 0.941 (.000) 0.941 (.001)
ranolazine 3 0.812 (.024) 0.819 (.018) 0.830 (.010) 0.770 (.005) 0.730 (.011)
scaffold_hop 5 0.531 (.016) 0.615 (.100) 0.568 (.017) 0.551 (.024) 0.531 (.010)
sitagliptin 3 0.501 (.081) 0.634 (.039) 0.677 (.055) 0.470 (.041) 0.402 (.024)
thiothixene 6 0.504 (.033) 0.583 (.034) 0.544 (.067) 0.544 (.026) 0.515 (.028)
troglitazone 4 0.437 (.067) 0.511 (.054) 0.487 (.024) 0.458 (.018) 0.420 (.031)
valsartan_smarts 2 0.158 (.317) 0.135 (.271) 0.000 (.000) 0.182 (.363) 0.119 (.238)
zaleplon 4 0.504 (.022) 0.552 (.033) 0.514 (.033) 0.533 (.009) 0.492 (.021)
Average 3.6 (1.3) 0.662 (.221) 0.705 (.219) 0.682 (.239) 0.660 (.213) 0.624 (.215)
Rank by avg. score 3 1 2 4 5
Table 14: Comparison of GP-MoLFormer-Sim+GA to main results in [kim2024geneticguidedgflownetssampleefficient] - Page 2 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task GP BO Fragment GFN Fragment GFN-AL
albuterol 0.902 (.011) 0.382 (.010) 0.459 (.028)
amlodipine 0.579 (.035) 0.428 (.002) 0.437 (.007)
celecoxib 0.746 (.025) 0.263 (.009) 0.326 (.008)
deco_hop 0.615 (.009) 0.582 (.001) 0.587 (.002)
DRD2 0.941 (.017) 0.480 (.075) 0.601 (.055)
fexofenadine 0.726 (.004) 0.689 (.003) 0.700 (.005)
GSK3 0.861 (.027) 0.589 (.009) 0.666 (.006)
isomers_c7 0.883 (.040) 0.791 (.024) 0.468 (.211)
isomers_c9 0.805 (.007) 0.576 (.021) 0.199 (.199)
JNK3 0.611 (.080) 0.359 (.009) 0.442 (.017)
median1 0.298 (.016) 0.192 (.003) 0.207 (.003)
median2 0.296 (.011) 0.174 (.002) 0.181 (.002)
mestranol 0.631 (.093) 0.291 (.005) 0.332 (.012)
osimertinib 0.788 (.005) 0.787 (.002) 0.785 (.003)
perindopril 0.494 (.006) 0.423 (.006) 0.434 (.006)
QED 0.937 (.002) 0.904 (.002) 0.917 (.002)
ranolazine 0.741 (.010) 0.626 (.005) 0.660 (.004)
scaffold_hop 0.535 (.007) 0.461 (.002) 0.464 (.003)
sitagliptin 0.461 (.057) 0.180 (.012) 0.217 (.022)
thiothixene 0.544 (.038) 0.261 (.004) 0.292 (.009)
troglitazone 0.404 (.025) 0.183 (.001) 0.190 (.002)
valsartan_smarts 0.000 (.000) 0.000 (.000) 0.000 (.000)
zaleplon 0.466 (.025) 0.308 (.027) 0.353 (.024)
Average 0.620 (.232) 0.432 (.227) 0.431 (.219)
Rank by avg. score 6 7 8
Table 15: Comparison of GP-MoLFormer-Sim+GA to main results in [wang2024efficient]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
Task Our Our REINVENT Augmented Graph GP BO MOLLEO MOLLEO MOLLEO
Rank AUC Memory GA GP BO (MolSTM) (BioT5) (GPT-4)
albuterol 8 0.824 (.071) 0.896 (.008) 0.918 (.026) 0.874 (.020) 0.902 (.019) 0.929 (.005) 0.968 (.003) 0.985 (.024)
amlodipine 4 0.680 (.064) 0.642 (.044) 0.686 (.046) 0.625 (.040) 0.552 (.025) 0.674 (.018) 0.776 (.038) 0.773 (.037)
celecoxib 5 0.716 (.067) 0.716 (.084) 0.784 (.011) 0.582 (.057) 0.728 (.048) 0.594 (.105) 0.508 (.017) 0.864 (.034)
deco_hop 3 0.710 (.058) 0.666 (.044) 0.688 (.060) 0.619 (.004) 0.629 (.018) 0.613 (.016) 0.827 (.093) 0.942 (.013)
DRD2 6 0.956 (.010) 0.945 (.007) 0.962 (.005) 0.964 (.012) 0.923 (.017) 0.975 (.003) 0.981 (.002) 0.968 (.012)
fexofenadine 2 0.798 (.028) 0.769 (.009) 0.686 (.010) 0.779 (.025) 0.745 (.009) 0.789 (.016) 0.773 (.017) 0.847 (.018)
GSK3 2 0.896 (.035) 0.865 (.043) 0.889 (.027) 0.788 (.070) 0.851 (.041) 0.898 (.041) 0.889 (.015) 0.863 (.047)
isomers_c7 5 0.932 (.011) 0.842 (.029) 0.954 (.033) 0.949 (.036) 0.662 (.071) 0.948 (.036) 0.928 (.038) 0.984 (.008)
isomers_c9 4 0.864 (.016) 0.642 (.054) 0.830 (.016) 0.719 (.047) 0.469 (.180) 0.871 (.039) 0.873 (.019) 0.874 (.053)
JNK3 1 0.806 (.087) 0.783 (.023) 0.773 (.073) 0.553 (.136) 0.564 (.155) 0.643 (.226) 0.728 (.079) 0.790 (.027)
median1 3 0.340 (.034) 0.372 (.015) 0.335 (.012) 0.287 (.008) 0.325 (.012) 0.298 (.019) 0.338 (.033) 0.352 (.024)
median2 6 0.255 (.031) 0.294 (.006) 0.290 (.006) 0.229 (.017) 0.308 (.034) 0.251 (.031) 0.259 (.019) 0.275 (.045)
mestranol 4 0.658 (.118) 0.618 (.048) 0.764 (.035) 0.579 (.022) 0.627 (.089) 0.596 (.018) 0.717 (.104) 0.972 (.009)
osimertinib 5 0.819 (.004) 0.834 (.046) 0.856 (.013) 0.808 (.012) 0.762 (.029) 0.823 (.007) 0.817 (.016) 0.835 (.024)
perindopril 4 0.584 (.042) 0.537 (.016) 0.598 (.008) 0.538 (.009) 0.493 (.011) 0.554 (.037) 0.738 (.016) 0.600 (.031)
QED 4 0.940 (.001) 0.941 (.000) 0.941 (.000) 0.940 (.000) 0.937 (.000) 0.937 (.002) 0.937 (.002) 0.948 (.000)
ranolazine 1 0.812 (.024) 0.760 (.009) 0.802 (.003) 0.728 (.012) 0.735 (.013) 0.725 (.040) 0.749 (.012) 0.769 (.022)
scaffold_hop 6 0.531 (.016) 0.560 (.019) 0.565 (.008) 0.517 (.007) 0.548 (.019) 0.527 (.019) 0.559 (.102) 0.971 (.004)
sitagliptin 4 0.501 (.081) 0.021 (.003) 0.479 (.039) 0.433 (.075) 0.186 (.055) 0.548 (.065) 0.506 (.100) 0.584 (.067)
thiothixene 7 0.504 (.033) 0.534 (.013) 0.562 (.028) 0.479 (.025) 0.559 (.027) 0.508 (.035) 0.696 (.081) 0.727 (.052)
troglitazone 4 0.437 (.067) 0.452 (.048) 0.556 (.052) 0.377 (.010) 0.405 (.007) 0.381 (.025) 0.390 (.044) 0.562 (.019)
valsartan 2 0.158 (.317) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.000 (.000) 0.867 (.092)
zaleplon 2 0.504 (.022) 0.347 (.049) 0.438 (.082) 0.456 (.007) 0.272 (.026) 0.475 (.018) 0.465 (.026) 0.510 (.031)
Average 4.0 (1.8) 0.662 (.221) 0.610 (.260) 0.668 (.237) 0.601 (.239) 0.573 (.239) 0.633 (.245) 0.671 (.248) 0.777 (.200)
Rank by 4 6 3 7 8 5 2 1
avg. score
Table 16: The mean ($\pm$ standard deviation) Synthetic Accessibility scores (SA) of top 100 molecules generated for each task using GP-MoLFormer-Sim+GA. The mean values are less than the threshold of 6 suggested in [Ertl_Schuffenhauer_2009], indicating that the molecules are easily synthesizable.
Task SA
albuterol 4.235 (0.433)
amlodipine 3.453 (0.101)
celecoxib 2.422 (0.322)
deco_hop 3.831 (0.541)
DRD2 4.004 (0.401)
fexofenadine 5.285 (0.320)
GSK3 3.230 (0.425)
isomers_c7 2.834 (0.795)
isomers_c9 4.301 (0.226)
JNK3 (zincseed)+RFF768+XO+DIV 3.114 (0.371)
median1 (zincseed)+RFF768+XO+DIV 3.587 (0.041)
median2 (zincseed)+RFF768+XO+DIV 2.882 (0.353)
mestranol 3.836 (0.203)
osimertimib 4.031 (0.155)
perindopril 3.791 (0.289)
QED 2.417 (0.314)
ranolazine 4.040 (0.292)
scaffold_hop 3.621 (0.273)
sitagliptin 3.967 (0.460)
thiothixene 2.546 (0.360)
troglitazone 3.329 (0.228)
valsartan 3.066 (1.031)
zaleplon 3.461 (0.501)
B.4 Extended Results — Trypsin inhibitors
Algorithm 1 anticipates the possibility of guidance by multiple targets at the same time (see Line 19 of Algorithm 1 in Section 2.2). In order to maintain focus on high scoring candidates in the context of the GA-based optimization, we operated the guidance mechanism in a single-target mode in main experiments. However, the guidance mechanism via GP-MoLFormer-Sim operates equally well in multi-guide mode. This is exemplified visually in Figure 5 using the t-SNE projection. The top left plot shows single-guide generations for the five target compounds for reference. The top right and bottom plots show generations guided by two of the compounds simultaneously (NAPAMP+UK-156406 and Efegatran+UK-156406, respectively). In the multi-guide cases, a clear trend in the direction of the combined targets on the t-SNE plot can be observed indicating guidance efficacy. Such mode can be of practical use in applications involving multi-objective or group-wise molecular optimization.
Figure 5: Examples of guided generation operating in single-guide (top left) and multi-guide mode (top right, and bottom) using selected combinations of Trypsin inhibitors.
Appendix C Hyperparameter Settings
Typical values for essential hyperparameters are listed in Table 17.
Table 17: GP-MoLFormer-Sim+GA hyperparameters used.
Parameter Type Description Value Comment
Multi-guide generation Boolean Each guide considered a separate target True Otherwise average of guides taken
Number of guides Int Best-score candidates 3 May be subject to pruning due to low Oracle value. Diversity guides are additional.
Guide pruning percentage Float Guides below this % of the top guide are pruned 75.0
Post-gen pruning $K$ Integer Offspring’s Tanimoto to top-1 must be larger than that of the $K$-th best candidate before being sent to the Oracle 10 This is an exploitative step. Exploration arrangements are done separately
Exploration candidates Int Add up to this number of exploration candidates per generation 40
Exploration method String Method to select exploration candidates: “Random” or “Crossover” Varies In “Crossover” mode, best guides serve as parents to create novel candidates
Diversity guides Int number of diversity guides to add in each guided run 1 This is in addition to parameter “Number of guides” above
Guidance strength $\alpha$ and temperature $\tau$ Float See Algorithm 1 0.4, 0.25
Exploitation trigger Float Change schedule if the top-1 Oracle value exceeds this value 0.95 This usually implies the target has been hit and the remaining 9 of 10 candidates should now get as close as possible to the top-1. Strategy is switched from exploration to exploitation
Exploitation $\alpha,\tau$ Float Values when exploitation mode is active 0.4, 0.15 Guiding at low temperature produces candidate very close to the guide
Stop after no change Int With no progress after this many generations, quit 500
Maximum generation size Int At any generation cap number of candidates sent to the Oracle 120
Generation size per guide Int 20
RFF Dimension Int Number of RFF features 768
RFF (Entropy) Temperature Float Corresponds to variance in the Gaussian kernel aproximation 0.008
GP-MoLFormer embedding dimensionality Float 768 See [ross2025gpmolformerfoundationmodelmolecular]